OpenStudy (anonymous):

Suppose the point (8,12) is on the graph of y=f(X). Find a point on the graph of each function. a) y=1/4f(x) b) y=4f(x)

OpenStudy (aum):

y = f(x) (8, 12) is a point on the graph. So 12 = f(8) ---- (1) a) y = 1/4 * f(x) when x = 8, y = 1/4 * f(x) = 1/4 * f(8) From (1): f(8) = 12. Therefore, 1/4 * f(8) = 1/4 * 12 = 3 (8,3) is a point on this graph. b) y = 4f(x) when x = 8, y = 4 * f(x) = 4 * f(8) = 4 * 12 = 48 (8, 48) is a point on this graph.